OpenStudy (anonymous):

A tomato of mass 0.18 kg is dropped from a tall bridge. If the tomato has a speed of 11 m/s just before it hits the ground, what is the kinetic energy of the tomato?

OpenStudy (anonymous):

K.E=1/2 mv^2 K.E=1/2 (0.18kg)(11m/s)^2